OpenStudy (anonymous):

4z^2-16+15/2z^2+z-15 simplifed form? 4z^2-16+15/2z^2+z-15 simplifed form? @Mathematics

OpenStudy (anonymous):

try to solve it by factorizing numerator and denominator first

OpenStudy (anonymous):

yasmeen i feel i sud tell u full procedure

OpenStudy (anonymous):

u think u can explain i t actually that would be great?

OpenStudy (anonymous):

is their z in middle term of the numerator

OpenStudy (anonymous):

4z^2 - 16 + 15 or, 4z^2 -16z + 15

OpenStudy (anonymous):

OHHH i wrote it wrong....... 4z^2-16+15/2z^2+z-15

OpenStudy (anonymous):

I did it again!

OpenStudy (anonymous):

Oh no its right

OpenStudy (anonymous):

\[\frac{4z^2 -16 + 15}{2z^2 + z-15}\] \[\frac{4z^2 -1}{2z^2 + 6z-5z-15}\] \[\frac{(2z)^2 -1}{2z(z + 3)-5(z+3)}\] \[\frac{(2z+1)(2z-1)}{(z + 3)(2z-5)}\]

OpenStudy (anonymous):

if that was right so it is right

OpenStudy (anonymous):

see i have choices and all the chioces look somthing like this 2z-3/z+3

OpenStudy (anonymous):

2z-3/z+3..... 2z+3/z+3... 2z+3/z-3.... 2z-3/z-3..... these are my choices

OpenStudy (anonymous):

\[\frac{4z^2 -10z -6z + 15 }{2z^2+6z-5z-15}= \frac{(2z-5)(2z-3)}{(2z-5)(z+3)} =\frac{2z-3}{z+3}\]

OpenStudy (anonymous):

thanks and good explaintion

OpenStudy (anonymous):

u think u can help me with a muliplication like this problem?

OpenStudy (anonymous):

if u want to learn then only otherwise how can any1......think of it ????

OpenStudy (anonymous):

1/5xy times 10x^2/7y

OpenStudy (anonymous):

\[\frac{2x}{7y}\]

OpenStudy (anonymous):

i was thinking it was 7/50x^3

OpenStudy (anonymous):

xy is not in the denominator

OpenStudy (anonymous):

mmmm ok

OpenStudy (anonymous):

can u walk me through this one because i dont think im understanding/// 12y4/88x^3 times 6x^8/20y^6

OpenStudy (anonymous):

\[\frac{xy}{5} \times \frac{10x^2}{7y} = \frac{2x^3}{7}\]

OpenStudy (anonymous):

oh so its like cross mulitplying?

OpenStudy (anonymous):

its not 88! its just 8x^3

OpenStudy (anonymous):

\[\frac{12y^4}{8x^3} \times \frac{6x^8}{20y^6} = \frac{9x^5}{20y^2}\]

OpenStudy (anonymous):

thanks!\